OpenStudy (anonymous):

what is the system of equations for this ordered pair 5y-3x=16 4x-3y=-3

OpenStudy (anonymous):

5y-3x=16 ----(1) 4x-3y=-3 ----(2) multiplying eq1by 4 and eq 2 by 3 and adding them we find: 20y -12x=64 12x-9y=-9 ______________ 11y=55 i.e. y= 55/11 i.e. y=5 substituting this value of y in eq (1) we find 10-3x=16 i.e. -3x= 16-10 i.e. -3x= 6 i.e. x= -2 thus (x,y) = (-2, 5) is the ordered pair for the given system of equations